# Context-Driven Knowledge Graph Completion with Semantic-Aware Relational Message Passing

### v1

@@ 2.2 Graph Neural Networks
Graph Neural Networks (GNNs) [14, 15, 16, 17] have been introduced to KGs due to their strong modeling capabilities on graph-structured data. Methods like R-GCN [17] and CompGCN [16] update entity representations by aggregating information from neighbors. However, this indiscriminate aggregation can introduce excessive noise and may lead to the over-smoothing problem, where entity representations become indistinguishable after multiple message-passing layers. While some subsequent methods (e.g., conceptually similar to RED-GNN [14]) attempt to mitigate this using soft attention mechanisms to select edges, the risk of noise and information dilution persists as all edges (even with low weights) can still contribute to the final representation, potentially obscuring crucial semantic cues around a target triplet $(h,r,t)$.
@@ 3.1 Preliminaries
A knowledge graph is represented as $K=(\mathcal{V},\mathcal{E},\mathcal{F})$, where $\mathcal{V}$ is the set of entities, $\mathcal{E}$ is the set of relations, and =$\mathcal{F}=\{(h,r,t)\mid h,t\in\mathcal{V},r\in\mathcal{E}\}$ is the set of factual triplets. Given a query entity $h$, a query relation $r$, and an unknown answer entity $t$, the task of knowledge graph completion is to predict $t$ for the query $(h,r,?)$. Typically, any entity in $\mathcal{V}$ is a candidate for $t$.
@@
3.2.1 Semantic-Aware Top-K Edge Selection Strategy
A cornerstone of our enhanced message passing paradigm is the semantic-aware Top-K edge selection strategy. This strategy mitigates noise and information dilution inherent in indiscriminate aggregation. For a given node $v$ with its current representation $\mathbf{h}_{v}^{(l)}$, we assess the semantic relevance between $\mathbf{h}_{v}^{(l)}$ and each incident edge $e\in\mathcal{N}(v)$, where each edge is characterized by its current state representation $\mathbf{s}_{e}^{(l)}$. This evaluation involves projecting $\mathbf{h}_{v}^{(l)}$ and $\mathbf{s}_{e}^{(l)}$ into a shared latent semantic space via a learnable mapping function $f(\cdot)$. Within this space, their similarity score is computed as:
$\text{Score}(v,e)=\exp(-\|f(\mathbf{h}_{v}^{(l)})-f(\mathbf{s}_{e}^{(l)})\|^{2}/\tau),$ (3)
where $\|\cdot\|^{2}$ denotes the squared Euclidean distance, and $\tau$ is a temperature hyperparameter. Only the Top-K edges $\mathcal{N}_{K}(v)\subseteq\mathcal{N}(v)$ exhibiting the highest scores, thus deemed most semantically pertinent, are chosen for the subsequent message construction phase. This selective mechanism ensures that only the most relevant information contributes to the node’s updated representation.
3.2.2 Multi-Head Attention Aggregator.
Upon identifying the Top-K semantically relevant edges $\mathcal{N}_{K}(v)$, their information must be effectively integrated with the node’s own representation $\mathbf{h}_{v}^{(l)}$ to derive its updated representation $\mathbf{h}_{v}^{(l+1)}$. To this end, we employ a multi-head attention aggregator. The states of the selected Top-K edges, $\{\mathbf{s}_{e}^{(l)}\}_{e\in\mathcal{N}_{K}(v)}$, are first consolidated into a unified neighborhood representation, for instance, through mean aggregation:
$\mathbf{\bar{s}}_{\mathcal{N}_{K}(v)}^{(l)}=\text{Mean}\left(\{\mathbf{s}_{e}^{(l)}\}_{e\in\mathcal{N}_{K}(v)}\right).$ (4)
This consolidated neighborhood representation $\mathbf{\bar{s}}_{\mathcal{N}_{K}(v)}^{(l)}$ and the node’s current representation $\mathbf{h}_{v}^{(l)}$ are then processed by the multi-head attention aggregator. The aggregator transforms $\mathbf{h}_{v}^{(l)}$ and $\mathbf{\bar{s}}_{\mathcal{N}_{K}(v)}^{(l)}$, computes head-wise gating scores based on their interaction, applies these gates to the transformed $\mathbf{\bar{s}}_{\mathcal{N}_{K}(v)}^{(l)}$, and then combines this gated representation with the transformed $\mathbf{h}_{v}^{(l)}$ before a final linear projection. This procedure yields the semantically focused updated node representation $\mathbf{h}_{v}^{(l+1)}$:
$\mathbf{h}_{v}^{(l+1)}=\text{Multi-HeadAttention}\left(\mathbf{h}_{v}^{(l)},\mathbf{\bar{s}}_{\mathcal{N}_{K}(v)}^{(l)}\right).$ (5)
@@ 3.2.1 Semantic-Aware Top-K Edge Selection Strategy
where []delimited-[][$\cdot$][ ⋅ ] denotes concatenation. This edge message $\mathbf{m}_{e}^{(l+1)}$ is subsequently used to update the hidden state of edge $e$:
@@ 3.3 Training And Inference
Here, $[\mathbf{m}_{h},\mathbf{x}_{e},\mathbf{m}_{t}]$ denotes the concatenation of the head entity message, the relation representation, and the tail entity message.

### v2

@@ 2.2 Graph Neural Networks
Graph Neural Networks (GNNs) [14, 15, 16, 17, 48] have been introduced to KGs due to their strong modeling capabilities on graph-structured data. Methods like R-GCN [17] and CompGCN [16] update entity representations by aggregating information from neighbors. However, this indiscriminate aggregation can introduce excessive noise and may lead to the over-smoothing problem, where entity representations become indistinguishable after multiple message-passing layers. While some subsequent methods (e.g., conceptually similar to RED-GNN [14]) attempt to mitigate this using soft attention mechanisms to select edges, the risk of noise and information dilution persists as all edges (even with low weights) can still contribute to the final representation, potentially obscuring crucial semantic cues around a target triplet $(h,r,t)$.
@@ 3.1 Preliminaries
A knowledge graph is represented as $K=(\mathcal{V},\mathcal{E},\mathcal{F})$, where $\mathcal{V}$ is the set of entities, $\mathcal{E}$ is the set of relations, and $\mathcal{F}=\{(h,r,t)\mid h,t\in\mathcal{V},r\in\mathcal{E}\}$ is the set of factual triplets. Given a query entity $h$, a query relation $r$, and an unknown answer entity $t$, the task of knowledge graph completion is to predict $t$ for the query $(h,r,?)$. Typically, any entity in $\mathcal{V}$ is a candidate for $t$.
@@
The architecture of the semantic-aware relational message passing model. The left side illustrates the overall workflow: a central edge selects its Top-K neighbors based on semantic similarity, aggregates their information, and updates its state. The right side shows a detailed diagram of the multi-head attention aggregator, which processes the central edge’s state and the aggregated neighbor state to produce the final output.
@@
A cornerstone of our enhanced message passing paradigm is the semantic-aware Top-K edge selection strategy, which refines how a node’s representation is updated. The entire process begins by contextually enriching each edge $e_{v}\in\mathcal{N}(v)$ incident to a target node $v$. For each such edge, treated as a central edge, we update its state by selectively aggregating information from its own local neighborhood of edges, $\mathcal{N}(e_{v})$. This mitigates noise and information dilution inherent in indiscriminate aggregation. Specifically, for a central edge $e_{v}$ with its current state $\mathbf{s}_{e_{v}}^{(l)}$, we assess its semantic relevance to each neighboring edge $e_{n}\in\mathcal{N}(e_{v})$ by projecting their states into a shared latent semantic space via a learnable mapping function $f(\cdot)$. Within this space, their similarity score is computed as:
$\text{Score}(e_{v},e_{n})=\exp(-\|f(\mathbf{s}_{e_{v}}^{(l)})-f(\mathbf{s}_{e_{n}}^{(l)})\|^{2}/\tau),$ (3)
where $\|\cdot\|^{2}$ denotes the squared Euclidean distance, and $\tau$ is a temperature hyperparameter. Only the Top-K edges $\mathcal{N}_{K}(e_{v})\subseteq\mathcal{N}(e_{v})$ exhibiting the highest scores are chosen. The states of these selected edges, $\{\mathbf{s}_{e_{n}}^{(l)}\}_{e_{n}\in\mathcal{N}_{K}(e_{v})}$, are first consolidated into a unified neighborhood representation through mean aggregation:
$\mathbf{\bar{s}}_{\mathcal{N}_{K}(e_{v})}^{(l)}=\text{Mean}\left(\{\mathbf{s}_{e_{n}}^{(l)}\}_{e_{n}\in\mathcal{N}_{K}(e_{v})}\right).$ (4)
This aggregated representation and the central edge’s own state are then processed by a multi-head attention aggregator to produce a contextually-aware, updated edge state $\mathbf{s}_{e_{v}}^{(l+1)}$. After this enrichment procedure is performed for every edge incident to node $v$, we obtain a set of updated edge representations $\{\mathbf{s}_{e_{v}}^{(l+1)}\}_{e_{v}\in\mathcal{N}(v)}$. The final updated representation for the node, $\mathbf{h}_{v}^{(l+1)}$, is then derived by applying mean pooling to these enriched incident edge states. This approach defines the node’s identity purely through the contextualized information of its surrounding relational structure, without direct use of its previous state $\mathbf{h}_{v}^{(l)}$:
$\mathbf{h}_{v}^{(l+1)}=\text{Mean}\left(\{\mathbf{s}_{e_{v}}^{(l+1)}\}_{e_{v}\in\mathcal{N}(v)}\right).$ (5)
@@
where $\cdot$] denotes concatenation. This edge message $\mathbf{m}_{e}^{(l+1)}$ is subsequently used to update the hidden state of edge $e$:
@@ 3.3 Training And Inference
Here, $\mathbf{m}_{h},\mathbf{x}_{e},\mathbf{m}_{t}]$ denotes the concatenation of the head entity message, the relation representation, and the tail entity message.
@@ 4.3 Ablation Study: RQ2
A line chart showing model performance (MRR and Hits@1) for different Top-K values. Performance peaks at K equals 10 and then declines.
@@
A bar chart comparing performance for different context aggregators. The Multi-Head Attention aggregator achieves the highest scores for both MRR and Hits@1.
@@
A line chart showing model performance versus the number of hops. The performance is optimal at 2 hops and degrades with more hops.
@@ 5 Conclusion
6 Conflict of Interest Statement
The authors have no competing interests.